You are working on a bid to build two city parks a year for the next three years. This project requires the purchase of $273,000 of equipment that will be depreciated using straight-line depreciation to a zero book value over the three-year project life. Ignore bonus depreciation. The equipment can be sold at the end of the project for half of what you paid for it. You will also need $23,000 in net working capital for the duration of the project. The fixed costs will be $64,500 a year and the variable costs will be $101,000 per park. Your required rate of return is 16 percent and your tax rate is 21 percent.

a. What is the depreciation per year? $

b.  After-tax salvage value of equipment is:  $
Answer #1

a)

Depreciation per year = Cost of equipment / number of years

Depreciation per year = 273,000 / 3

Depreciation per year = $91,000

b)

Sale value = 273,000 / 2 =136,500

after tax salvage value = Market value - tax(market value - book value)

after tax salvage value = 136,500 - 0.21(136,500 - 0)

after tax salvage value = 136,500 - 28,665

after tax salvage value = $107,835